OpenStudy (anonymous):

A cone with a radius of 8 feet and a slant height of 17.....surface area?

OpenStudy (rsadhvika):

Surface area = \(\pi r^2 + \pi r l\) \(r\) = radius \(l\) = slant height

OpenStudy (souvik):

surface area=(pi*radius*slant height)+(pi*radius^2)

OpenStudy (anonymous):

thanks u2 lol i have 5 virtuals i have to finish

OpenStudy (rsadhvika):

you're doing great ! yw :)